three times the larger of two numbers is equal to four times the smaller. The sum of the numbers is 21 (12+9). Find the numbers.
Answer by checkley77(12569)

X+Y=21 OR X=21-Y
3Y=4X
3Y=4(21-Y)
3Y=84-4Y
3Y+4Y=84
7Y=84
Y=84/7
Y=12 THE LARGER NUMBER.
X+12=21
X=21-12
X=9 THE SMALLER NUMBER.
PROOF:
3*12=4*9
36=36